OpenStudy (anonymous):

Complete these ordered pairs for this equation. (0, ), (-1, ), (1, ) y = 2x + 1

OpenStudy (cwrw238):

just substitute for x ( = 0,-1,1) in the equation for x = 0 y = 2(0) + 1 = 1 so first ordered pair is (0, 1) repeat for x = -1 and 1

OpenStudy (anonymous):

what do u mean i dont getthis one help me out cuz i am stuck on this problem

OpenStudy (cwrw238):

the first value on the ordered pair is value of x the second number in the ordered pair is the value of y for the specified value of x so for the second ordered pair ( -1, ), x = -1 and to find y we plug in -1 into the equation y = 2x + 1 so for x = -1 , y = 2(-1) + 1 = -2 + 1 = -1 so second ordered pair =( x , y) = (-1, -1)

OpenStudy (anonymous):

i dont get you

OpenStudy (cwrw238):

we are given the x values and to get the y values ( that is the second number in the parentheses we put the x value instead of x in the equation y = 2x + 1 (0 , ) :- the x value is 0 We need to fill in the blank (which is the value of y when x = 0) y = 2x + 1 is the equation relating the values of x and y when x = 0 :- y = 2 times 0 + 1 = 0 + 1 = 1 so our ordered pair is (0 ,1)

OpenStudy (anonymous):

wirte all the pairs down that you have done